OpenStudy (anonymous):

15 micro farad and 20 micro farad are connected in series to a 600 V DC supply. i) find charge on each capacitor. ii) V across each resistor.

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

i Find Ceq by using 1/Ceq=1/C1+1/C2 Then Q=VCeq The charge on both the capacitors is same i.e. Q ii V1=Q/C1 V2=Q/C2

OpenStudy (anonymous):

galat hai

OpenStudy (anonymous):

the capacitors is serie have the same charge, then :\[q _{1}=15\mu F*V _{1}\] and \[q _{2}=20\mu F*V _{2}-->q _{1}=q _{2}\].. also, \[V _{1}+V _{2}=600V\].. you have the nest system of equations: \[15\mu F \times V_{1}=20\mu F \times V_{2}--> and-->V_{2}=600-V_{1}\] resolve and you will get \[V_{1} =600\times(\frac{ 20 }{35 })--> and-->V_{2}=600-600(\frac{ 20 }{ ? })\] :D